OpenStudy (rianiscool):

need help asap will give a medal A plumber charges $50 to come to a house and $20 an hour for work time. If h represents the number of hours worked, which expression gives the total amount the plumber charges? A. 50h + 20 B. 20h + 50 C. 20(h + 50) D. + 50

OpenStudy (rianiscool):

is is c

OpenStudy (texaschic101):

yes it is

OpenStudy (chosenmatt):

b

OpenStudy (rianiscool):

why b

OpenStudy (texaschic101):

my bad...I meant B....I looked at the wrong one

OpenStudy (chosenmatt):

yay im right!

OpenStudy (rianiscool):

ok thank you

OpenStudy (texaschic101):

he charges 50 to come to your house.....he then charges 20 per hour 20h + 50 --- lets say he worked 10 hrs 20(10) + 50 = 200 + 50 = 250...it would cost 250 understand ?